OpenStudy (anonymous):

(1/1-sinx) + (1/1+sinx)=2sec^2x

OpenStudy (anonymous):

first you have to multiply by the common denominator

OpenStudy (anonymous):

Did it ??

OpenStudy (anonymous):

What do you mean

OpenStudy (anonymous):

i was asking if you have solve it cuz it took u too long to repy me ;)

OpenStudy (anonymous):

Im sorry im doing this on my phone its kinda hard to do. No I didn't solve it yet

OpenStudy (anonymous):

it`s easy though im sure u can solve it very easily ;)

OpenStudy (anonymous):

1/(1- sin x) + 1/(1 + sin x) = [ (1 + sin x) + (1 - sin x)]/(1 - sin^2 x) = =2/cos^2 x